If 10.0 mL of 2.50 M SrS (aq) is made to react with 25.0 mL of 1.50 M ZnSO4 (aq), how many moles of the excess reactant will be left after the reaction? Express your answer in three significant figures.


1
Expert's answer
2021-04-27T07:47:57-0400

ZnSO4+SrCl2 "\implies" ZnCl2+SrSO4

ZnSO4+SrS"\implies" ZnS+SrSO4

But SrS(aq)=Volume (10mL)

M(2.50M)


"Moles=\frac{V×M}{1000}"


"=\frac{10×2.5}{1000}" "=0.025moles"


ZnSO4(aq)="\frac{V×M}{1000}"


="\frac{25×1.5}{1000}=0.0375moles"


Therfore:

Moles of ZnS using mole ratio="0.0375moles"

Moles of SrSO4 using mole ratio =

"0.025moles"

Excess reactants ZnSO4 moles produced="(0.0375-0.025)moles"


"=0.0125moles"